Solve for " h "
A=1/2 bh
Thanks
: Solve for " h "
A=1/2 bh
Thanks

Answer by nerdybill(1052)

To solve for 'h', simply isolate it to one side of the equation -- doesn't matter which side.
.
A=1/2 bh
.
Begin by multiplying both sides by 2:
A=1/2 bh
2A = bh
.
Finally, divide both sides by 'b':
(2A)/b = h